# Towards Personalized Treatment Plan: Geometrical Model-Agnostic Approach to Counterfactual Explanations

### v1

@@ 6 Conclusion
In the current line of research for counterfactual explanations, much of the recent research in the space of counterfactual generation has focused on optimization (hyperparameter tuning, gradient descent) and the method of tuning the parameters of a loss function that reduces the $L_{2}$ norm between the original instance and newly generated counterfactual $\mathbf{x^{\prime}}$ where $n$ is the number of features. In DiCE’s case, they have both model-agnostic and gradient-based approaches that use a loss function to optimize the original counterfactual explanation. Their genetic algorithm, one of their three model-agnostic approaches, applies a loss function for optimizing an original set of counterfactuals for gradual improvement across multiple iterations (Mothilal et al., 2020). Mahajan et. al. [2019] applied both gradient descent and hyperparameter tuning. They searched for possible counterfactuals using “random search for $100$ iterations” to best fit the objective that they designed (Mahajan et al., 2019). With our work, we propose a model-agnostic method that produces CFEs in high-dimensional space based on the selection of decision boundary points that minimize the $L_{2}$ norm between the original instance and the boundary point. This method performs well compared to other current model-agnostic methods and optimization techniques with enough boundary points when focusing on the Euclidean metric.
@@ 7 Future Works
Finally, regarding questions of feasibility and the search for a personalized treatment plan, a method where we can identify mutable features from immutable ones could also provide directional improvement towards a “no disease” scenario. We can also utilize more inequalities as constraints rather than equalities and use similarity rather than an exact match. Given that finding a decision boundary point with the same categorical features will be unlikely (i.e., finding a person with the same age, same gender, same sex, etc.), we can use inequalities to help find a decision boundary point that is most similar to the original query point. To explain this in simpler terms, if an observational dataset consisted of persons with a known target label for a given disease diagnosis, we might not find a boundary point that equates in all categorical features of the original query (person), such as age, sex, gender, and number of times the person smoked each week. However, if we relax the problem with inequalities, we might find a similar boundary point that provides a case for directional improvement. In a similar sense, Mahajan et. al. [2019] used an inequality for the age of a person in the Adult Income dataset rather than a strict equality (Mahajan et al., 2019).

### v4

@@ 5 Conclusion
In the current line of research for counterfactual explanations, much of the recent research in the space of counterfactual generation has focused on optimization (hyperparameter tuning, gradient descent) and the method of tuning the parameters of a loss function that reduces the $L_{2}$ norm between the original instance and newly generated counterfactual $\mathbf{x^{\prime}}$ where $n$ is the number of features. In DiCE’s case, they have both model-agnostic and gradient-based approaches that use a loss function to optimize the original counterfactual explanation. Their genetic algorithm, one of their three model-agnostic approaches, applies a loss function for optimizing an original set of counterfactuals for gradual improvement across multiple iterations (mothilal2020dice). Mahajan et. al. [2019] applied both gradient descent and hyperparameter tuning. They searched for possible counterfactuals using “random search for $100$ iterations” to best fit the objective that they designed (mahajan2019preserving). With our work, we propose a model-agnostic method that produces CFEs in high-dimensional space based on the selection of decision boundary points that minimize the $L_{2}$ norm between the original instance and the boundary point. This method performs well compared to other current model-agnostic methods and optimization techniques with enough boundary points when focusing on the Euclidean metric.
@@ 6 Future Works
We can also apply a small shifting factor to the lower bound of the decision boundary points separated by some epsilon $\epsilon$. If we shift the lower bound of points (closer to the original instance), we can sample and choose the minimum of points from the shifts.
@@
Finally, regarding questions of feasibility and the search for a personalized treatment plan, a method where we can identify mutable features from immutable ones could also provide directional improvement towards a “no disease” scenario. We can also utilize more inequalities as constraints rather than equalities and use similarity rather than an exact match. Given that finding a decision boundary point with the same categorical features will be unlikely (i.e., finding a person with the same age, same gender, same sex, etc.), we can use inequalities to help find a decision boundary point that is most similar to the original query point. To explain this in simpler terms, if an observational dataset consisted of persons with a known target label for a given disease diagnosis, we might not find a boundary point that equates in all categorical features of the original query (person), such as age, sex, gender, and number of times the person smoked each week. However, if we relax the problem with inequalities, we might find a similar boundary point that provides a case for directional improvement. In a similar sense, Mahajan et. al. [2019] used an inequality for the age of a person in the Adult Income dataset rather than a strict equality (mahajan2019preserving).
@@
7 Appendix
7.1 Evaluation Metrics
In the results section, we share the final results from evaluating various methodologies, including our Segmented Sampling from Boundary Approximation approach (SSBA), Alibi’s gradient-based approach, and DiCE’s model-agnostic approaches. In this section, we present the design of the metrics used to evaluate our method in comparison to Alibi and DiCE’s approaches.
As shown in Table $1(b)$, we compute the distance column (last column) as an average over all randomly sampled points shown under the number of samples column.
In the unconstrained case, we compute the scalar distance as an average over the sampled points.
The formula that we use for the scalar in the last column (no constraints) is the one below:
$\frac{1}{|B|}*\Sigma_{(x,x^{\prime})\ \in\ B\ \subseteq\ X_{1}\ \times\ X^{\prime}\ }\ l_{2}(x,x^{\prime})$
This dataset $B$ is the sample of points from $X_{1}\ \times X^{\prime}$, where $X_{1}$ is our set of points that has the label $y=1$ in our dataset, and $X^{\prime}$ is the set of all possible generated counterfactuals. In simple terms, we compute the average distances over all data points that are labelled $1$ or “unhealthy” in the case of the toy dataset. We sample all points from $x\in X_{1}$ according to a uniform distribution, which means that $x\sim Unif(X_{1})$. $x^{\prime}\in X^{\prime}$ is the generated counterfactual constructed with $x\in X_{1}$. For computing the distance, we use $l_{2}$, which is the Euclidean metric on $\mathbf{R}^{n}$. The $l_{2}$ norm function computes the Euclidean distance between $x,x^{\prime}\in\mathbf{R}^{n}$, where $n$ is the dimension of the space.
In the constrained case, we compare the distance from a given generated counterfactual $x^{\prime}\in\Delta X^{\prime}$ to one of the decision boundary points generated with the SSBA method. We define all decision boundary points as points $d\in D$ where $D$ is our total set of generated decision boundary points. $\Delta X^{\prime}$ is the set of constrained counterfactuals where only the continuous features are constrained and categorical features are left unchanged.
[…]
$\frac{1}{|B|}*\Sigma_{(d^{*},x^{\prime})\ \in\ B\ \subseteq\ D\ \times\ \Delta X^{\prime}\ }\ l_{2}(d,x^{\prime})$
7.2 Metric Examples
Figure 5: This is an example of how we compute the distance values in the last column of Table $1(b)$ for the unconstrained case. We sample the distances in black where the points are labelled $y=1$ (or unhealthy in this case). We then average over these distances.
Figure 6: This is an example of how we compute the distance values in the last column of Table $1(b)$ for the constrained case. We sample the distances from the bounded counterfactual highlighted in green to the closest decision boundary point. We sample from the points labelled $y=1$ (or unhealthy in this case) and average the distances.
7.3 Outline of Proof for Bisection Method Search
 [^2] Vlassopoulos et. al. (2020) do not provide a proof for the bisection method search that they suggested, so we will provide a proof here.
Assumption: For this proof, we apply the following definition for a decision boundary point $d$. For a point $d$, we define that the $l_{2}$ distance between $l_{2}(v_{1},d)$ and $l_{2}(v_{2},d)$ as part of two different classes which are both correctly classified and labeled, namely $v_{1}\in X_{c_{1}}^{correct}\subset\mathbf{R}^{n}$ and $v_{2}\in X_{c_{2}}^{correct}\subset\mathbf{R}^{n}$, to be less than an sufficient epsilon $\epsilon^{*}$. We must also have a trained model $f$ which generates a decision boundary.
Claim: For a binary classification problem, the bisection method generates an nonempty set $D$ of discrete decision boundary points.
Proof.
The proof for the bisection method search algorithm is based on graph coloring, where we maintain a 2-coloring of a graph $G$ and reduce the $l_{2}$ distance among points of different classes using binary search.
Consider a 2-colored graph $G$ from the classes of points that were both correctly classified and labeled $X_{c_{1}}^{correct}$ and $X_{c_{2}}^{correct}$. Let $X_{c_{1}}^{correct}$ be the set of nodes whose coloring is red, and let $X_{c_{2}}^{correct}$ be the set of nodes whose coloring is blue. Let the edges of the graph $G$ be between nodes of different colors, and let every edge $(v_{1},v_{2})\in X_{c_{1}}^{correct}\times X_{c_{2}}^{correct}$ be an edge in the graph $G$.
Then consider an edge $(v_{1},v_{2})$ between two vertices $v_{1}\in X_{c_{1}}^{correct}$ and $v_{2}\in X_{c_{2}}^{correct}$. Take the midpoint of these two vertices, and we find a new point $\frac{v_{1}+v_{2}}{2}=x$ for which we do not know the label. With the trained model $f$, we apply the model to the point $x$ such that $f$ produces a label $f(x)$ of one of two colors, red or blue.
In the case that the label for $x$ is red, we replace the endpoint $v_{1}$ that was originally colored red with our new midpoint $x$ such that our new smaller edge becomes $(x,v_{2})$. This new edge maintains the coloring, and the new $l_{2}$ norm is now $l_{2}(x,v_{2})<l_{2}(v_{1},v_{2})$.
In the case that the label for $x$ is blue, we replace the endpoint $v_{2}$ that was originally blue with our new midpoint $x$ such that we have a smaller edge $l_{2}(v_{1},x)<l_{2}(v_{1},v_{2})$ while maintaining the coloring of $G$.
Without loss of generality, we can repeat this process for new midpoints such that for step $k$, a given edge $(v_{1}^{k+1},v_{2}^{k+1})$ will have half the norm of $(v_{1}^{k},v_{2}^{k})$. Because of this repeated halving of the interval, we know that $l_{2}(v_{1}^{k},v_{2}^{k})=\frac{l_{2}(v_{1},v_{2})}{2^{k}}$. We can choose $N$ such that $\frac{l_{2}(v_{1}^{1},v_{2}^{1})}{2^{N}}<\epsilon^{*}$. By algebra, we can rearrange to find that $log_{2}\frac{l_{2}(v_{1},v_{2})}{\epsilon^{*}}<N$. If we choose $N^{*}=\lfloor log_{2}\frac{l_{2}(v_{1},v_{2})}{\epsilon^{*}}\rfloor+1$ such that $N^{*}\geq N$, we have found an integer $N^{*}$ that ensures the two endpoints of the edge are separated by an $l_{2}$ norm less than $\epsilon^{*}$. After $N^{*}$ iterations, we take the midpoint $M$ of $v_{1}^{N^{*}}$ and $v_{2}^{N^{*}}$ such that $l_{2}(M,v_{1}^{N*})<\epsilon^{*}$ and $l_{2}(M,v_{2}^{N*})<\epsilon^{*}$. Thus, $M$ represents a decision boundary point because $v_{1}^{N*}$ and $v_{2}^{N*}$ are points of different colors separated by less than $\epsilon^{*}$. Doing this for all edges in $G$, we get a discrete set of boundary points $D$. ∎
7.4 Additional Properties of SSBA
7.4.1 Reasonable Feature Bounds on Decision Boundary Points
Assumptions: Same as above.
Claim: For a dataset $X$ where $X_{i}$ refers to all observations of a given feature $i$, we can guarantee that the SSBA method generates decision boundary points in a “reasonable” range, which we define to be
$d_{i}\in[\min(X_{i}),\max(X_{i})]\quad\forall i\in\{1,\dots,n\},$
for a decision boundary point
$d=(d_{1},d_{2},\dots,d_{n}),\ d_{i}\in\mathbb{R},\ d\in\mathbb{R}^{n}.$
Proof.
Let $\min(X_{i})$ and $\max(X_{i})$ be the minimum and maximum values of feature $i$ across the dataset, for each $i\in\{1,\dots,n\}$.
Take a decision boundary point $d=(d_{1},\cdots,d_{n})$ generated with the SSBA approach. By construction, after a sufficient number of iterations, we have
$d_{i}=\alpha x_{i}+(1-\alpha)y_{i},\quad\text{for some }\alpha\in[0,1],$
where $x_{i},y_{i}\in X_{i}$.
Since $x_{i},y_{i}\in[\min(X_{i}),\max(X_{i})]$, it follows that
$\min(X_{i})\leq x_{i}\leq\max(X_{i})\quad\text{and}\quad\min(X_{i})\leq y_{i}\leq\max(X_{i}).$
Therefore,
$d_{i}=\alpha x_{i}+(1-\alpha)y_{i}\leq\alpha\max(X_{i})+(1-\alpha)\max(X_{i})=\max(X_{i}),$
$d_{i}=\alpha x_{i}+(1-\alpha)y_{i}\geq\alpha\min(X_{i})+(1-\alpha)\min(X_{i})=\min(X_{i}).$
Thus,
$\min(X_{i})\leq d_{i}\leq\max(X_{i}).$
Since this holds for all $i\in\{1,\dots,n\}$, we conclude that
$d_{i}\in[\min(X_{i}),\max(X_{i})]\quad\forall i\in\{1,\dots,n\}.$
7.4.2 Generating More Decision Boundary Points Than Number of Data Points in the dataset
Due to the $N^{2}$ nature of the algorithm, our method is generally capable of generating more boundary points than there are data points in our dataset. We can generally ensure that we can generate a number of data points such that $|D|>>N$ where $|D|$ is the cardinality or the size of our decision boundary set and $N$ is the number of total data points in our original dataset. The only exception to this would be a heavily skewed dataset like 1 dataset in one class and the N1N-1 data points in the other class. Then we would generate only N1N-1 boundary points. But, for a balanced dataset, we will have this property. This becomes especially true for balanced datasets with large $N$ such that $N^{2}$ nature for the algorithm takes over and enables a much greater generation of discrete boundary points.